# refrences.bib

### v1

@@ 3 Resolution of the feasible solution set for Problem (1)
Consider the problem stated in Example 1. According to Definition 5 and Table 4 (in Section 6), we obtain $\mathscr{J}_{1}=\{1,2,5\}$, $\mathscr{J}_{2}=\{1,8\}$, $\mathscr{J}_{3}=\{6,7,8\}$, $\mathscr{J}_{4}=\{2,3,4\}$, $\mathscr{J}_{5}=\{1,3,4,5\}$, $\mathscr{J}_{6}=\{8,9\}$, $\mathscr{J}_{7}=\{3,8,10\}$, $\mathscr{J}_{8}=\{9\}$, $\mathscr{J}_{9}=\{3,10\}$ and $\mathscr{J}_{10}=\{6,8\}$. Hence, according to Remark 5, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{I}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 3\times 4\times 2\times 3\times 1\times 2\times 2=5184$. Now, noting Definition 6 and Corollary 6, consider three functions $e_{1}=[5,1,6,3,5,8,10,9,10,6]$, $e_{2}=[1,8,6,2,3,8,10,9,10,6]$ and $e_{3}=[1,8,6,2,1,8,10,9,10,6]$ from $\mathscr{I}$ to $\bigcup_{i\in\mathscr{I}}\mathscr{J}_{i}$ so that $e_{p}(i)\in\mathscr{J}_{i}$ for each $i\in\mathscr{I}$ and each $p\in\{1,2,3\}$. For function $e_{1}$, we have $e_{1}(1)=e_{1}(5)=5$; $\mathscr{I}_{5}\left(e_{1},5\right)=\{1\}$ (from Part (II) of Definition 6); $S_{15}^{\prime}=\{0.6\}$ and $S_{55}^{\prime}=\{0.3\}$ (from Table 4). Therefore, $\mathscr{I}_{5}\left(e_{1},5\right)\neq\varnothing$ and $S_{55}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{5}\left(e_{1},5\right)}S_{k5}^{\prime}\right)=S_{55}^{\prime}\cap S_{15}^{\prime}=\varnothing$, which implies $5\notin\mathscr{J}_{5}\left(e_{1}\right)$ (Definition 6, Part (III)). So, since $e_{1}(5)=5\notin\mathscr{J}_{5}\left(e_{1}\right)$, from Definition 6 it follows that $e_{1}$ is not an admissible function. However, both $e_{2}$ and $e_{3}$ are admissible. For instance, for admissible function $e_{2}$ we have $\mathscr{I}_{1}\left(e_{2}\right)=\{1\}$, $\mathscr{I}_{2}\left(e_{2}\right)=\{4\}$, $\mathscr{I}_{3}\left(e_{2}\right)=\{5\}$, $\mathscr{I}_{4}\left(e_{2}\right)=\varnothing$, $\mathscr{J}_{5}\left(e_{2}\right)=\varnothing$, $\mathscr{I}_{6}\left(e_{2}\right)=\{3,10\}$, $\mathscr{I}_{7}\left(e_{2}\right)=\varnothing$, $\mathscr{I}_{8}\left(e_{2}\right)=\{2,6\}$, $\mathscr{I}_{9}\left(e_{2}\right)=\{8\}$ and $\mathscr{I}_{10}\left(e_{2}\right)=\{7,9\}$. So, by noting Table 4, it follows that $\bigcap_{i\in\mathscr{I}_{1}\left(e_{2}\right)}S_{i1}^{\prime}=S_{11}^{\prime}=\{0.4\}$, $\bigcap_{i\in\mathscr{I}_{2}\left(e_{2}\right)}S_{i2}^{\prime}=S_{42}^{\prime}=\{0.64\}$, $\bigcap_{i\in\mathscr{I}_{3}\left(e_{2}\right)}S_{i3}^{\prime}=S_{53}^{\prime}=\{0.6\}$, $\bigcap_{i\in\mathscr{I}_{6}\left(e_{2}\right)}S_{i6}^{\prime}=S_{36}^{\prime}\cap S_{10,6}^{\prime}=\{0,1\}$, $\bigcap_{i\in\mathscr{I}_{8}\left(e_{2}\right)}S_{i8}^{\prime}=S_{28}^{\prime}\cap S_{68}^{\prime}=\{0.2\}$, $\bigcap_{i\in\mathscr{I}_{9}\left(e_{2}\right)}S_{i9}^{\prime}=S_{89}^{\prime}=\{0.8\}$ and $\bigcap_{i\in\mathscr{I}_{10}\left(e_{2}\right)}S_{i,10}^{\prime}=S_{7,10}^{\prime}\cap S_{9,10}^{\prime}=\{0.6\}$.
@@
Furthermore, from Table 3, $I_{4}=[0,1]$, $I_{5}=[0.3,0.6]$ and $I_{7}=[0,0.55]$. Hence, from Definition 7, $S\left(e_{2}\right)$ is calculated as the Cartesian product $S\left(e_{2}\right)=\{0.4\}\times\{0.64\}\times\{0.6\}\times[0,1]\times[0.3,0.6]\times\{0,1\}\times[0,0.55]\times\{0.2\}\times\{0.8\}\times\{0.6\}$.
@@ 4 Optimal solutions of Problem (1)
    Clearly, $S^{*}\subseteq S\left(A^{+},A^{-},b\right)$. Also, from (11) and (15), it follows that if $x\left(e_{0}\right)\in F$ (for some $e_{0}\in E$ ), then $x\left(e_{0}\right)\in S\left(e_{0}\right)\subseteq\bigcup_{e\in E}S(e)$. So, by Theorem 3 , it is concluded that $F\subseteq S\left(A^{+},A^{-},b\right)$. To complete the proof, it is sufficient to show that any solution $x\in S\left(A^{+},A^{-},b\right)-F$ cannot be an optimal solution. For this purpose, let $x\in S\left(A^{+},A^{-},b\right)-F$. So, from Theorem 3, there exists at least one $e_{0}\in E$ such that $x\in S\left(e_{0}\right)$. Now, by (11) and (15), we have $x\left(e_{0}\right)_{j}\leq x_{j}$, $\forall j\in\mathscr{J}$; and since $x\notin F$, at least one of the preceding inequalities must be strict. Without loss of generality, assume that $x\left(e_{0}\right)_{j_{0}}<x_{j_{0}}$. Hence, $c_{j_{0}}x\left(e_{0}\right)_{j_{0}}+\sum_{j\in\mathscr{J}-\left\{j_{0}\right\}}c_{j}x\left(e_{0}\right)_{j}<c_{j_{0}}x_{j_{0}}+\sum_{j\in\mathscr{J}-\left\{j_{0}\right\}}c_{j}x_{j}$, which implies $x\notin S^{*}$. ∎
Theorem 4 also provides a necessary optimality condition in the sense that if $x^{*}$ is an optimal solution to Problem (1), then it must belong to $F$. In other words, each solution $x(e)\in F$ is an optimal candidate solution.
@@ 5.1 Simplification techniques for finding the feasible solutions set
Under the assumptions of Lemma 8, the $i_{0}$’th equation is a redundant constraint. Lemma 8 additionally provides a further idea that is concerned with the case in which certain $x_{j}$’s can be calculated straight away without solving the problem, but just by reflecting on the special characteristics of the problem. Due to this, in such cases, any parts involved in these $x_{j}$’s can be excluded. These considerations have led us to come up with the following reduction rule.
Corollary 8.
@@ 5.2 Simplification techniques for finding optimal solutions
The techniques described in this subsection are based on Theorem 4. These techniques reduce the size of the problem in such a way that either the set $F$ does not change or the search domain is reduced to a subset of $F$ that contains optimal solutions. In fact, by using these methods, some parts of the feasible region that do not contain optimal solutions may be removed. Therefore, these methods cannot be used to find the feasible region and are only suitable for finding optimal solutions.
Lemma 11.
@@
    It is sufficient to prove $x_{j_{0}}^{*}=L_{j_{0}}$ for any optimal solution $x^{*}$. But, since $x^{*}\in S^{*}$, Theorem 4 implies that $x^{*}\in F$; that is, $x^{*}=x(e)$ for some $e\in E$. Now the result follows from Lemma 12. ∎
Lemma 13.
Suppose that the simplification technique stated in Corollary 11 has been applied and there exist $j_{1},j_{2}\in\mathscr{J}$ such that $\mathscr{I}_{j_{1}}\neq\varnothing$ and $\mathscr{I}_{j_{1}}\subseteq\mathscr{I}_{j_{2}}$. Also, define $F_{0}=\left\{x(e):e\in E_{0}\right\}$ and $E_{0}=\left\{e\in E:\mathscr{I}_{j_{1}}(e)=\varnothing\right\}$.
@@
According to Lemma 13, we can reduce our search to set $F_{0}$ instead of set $F$. To achieve this, we neglect all the admissible functions $e\in E-E_{0}$ by removing each $i\in\mathscr{I}_{j_{1}}$ before selecting the vectors $e$ to construct solutions $x(e)$. However, this is equivalent to assigning $x_{j_{2}}^{*}=L_{j_{2}}$ ($x_{j_{2}}^{*}=U_{j_{2}}$ in part (b) of Lemma 13) and deleting the $j_{1}$’th column from the problem. The following reduction rule summarizes the previous discussion.
Corollary 13.
Suppose that the simplification technique stated in Corollary 11 has been applied.
@@ 5.3 Simplification strategies for finding optimal solutions
Definition 9.
Suppose that the simplification technique stated in Corollary 11 has been applied. An admissible function $e$ is said to be a modified function if $e(i)\in\bar{\mathscr{J}}_{i}(e)$ , $\forall i\in\mathscr{I}$, where
@@
Example 3.
Consider admissible functions $e_{2}=[1,8,6,2,3,8,10,9,10,6]$ and $e_{3}=[1,8,6,2,1,8,10,9,10,6]$ stated in Example 2. As calculated in Example 2, $\mathscr{J}_{5}=\{1,3,4,5\}$. Also, for admissible function $e_{2}$, we have $e_{2}(1)=1$, $e_{2}(2)=8$, $e_{2}(3)=6$ and $e_{2}(4)=2$. Furthermore, Part (II) of Definition 6 indicates that $\mathscr{I}_{1}\left(e_{2},5\right)=\{1\}$ and $\mathscr{I}_{3}\left(e_{2},5\right)=\mathscr{I}_{4}\left(e_{2},5\right)=\mathscr{I}_{5}\left(e_{2},5\right)=\varnothing$. On the other hand, from Table 4, it follows that $S_{51}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{1}\left(e_{2},5\right)}S_{k1}^{\prime}\right)=S_{51}^{\prime}\cap S_{11}^{\prime}=\{0.4\}$. Therefore, from Part (III) of Definition 6, we obtain $\mathscr{J}_{5}\left(e_{2}\right)=\{1,3,4,5\}$, which implies $\mathscr{J}_{5}\left(e_{2}\right)\cap\left\{e_{2}(1),e_{2}(2),e_{2}(3),e_{2}(4)\right\}=\{1,3,4,5\}\cap\{1,8,6,2\}=\{1\}\neq\varnothing$. So, $\bar{\mathscr{J}}_{5}\left(e_{2}\right)=\{1\}\quad$ and $e_{2}(5)=3\notin\bar{\mathscr{J}}_{5}\left(e_{2}\right)$, that means $e_{2}$ is not a modified function (Definition 9). However, function $e_{3}$ is an example of an admissible function that is also a modified function.
Theorem 5.
Suppose that the simplification technique stated in Corollary 11 has been applied. If $S\left(A^{+},A^{-},b\right)\neq\varnothing$ and $S^{*}$ denotes the set of optimal solutions for Problem (1), then $S^{*}\subseteq\bar{F}\subseteq F$.
@@
    From Definition 9, it is clear that $\bar{F}\subseteq F$. To complete the proof, we shall show that for each $e\in E-\bar{E}$, there exists some $\bar{e}\in\bar{E}$ such that $\sum_{j\in\mathscr{J}}c_{j}x(\bar{e})_{j}\leq\sum_{j\in\mathscr{J}}c_{j}x(e)_{j}$. For this purpose, let $e\in E-\bar{E}$, =$\mathscr{I}^{\prime}=\left\{i_{1}^{\prime},\ldots,i_{k}^{\prime}\right\}$ and $\mathscr{J}^{\prime}=\left\{j_{1}^{\prime},\ldots,j_{k}^{\prime}\right\}$ where $\mathscr{J}_{i_{p}^{\prime}}(e)\bigcap\left\{e(1),\ldots,e\left(i_{p}^{\prime}-1\right)\right\}\neq\varnothing$ and $e(i_{p}^{\prime})=j_{p}^{\prime}\in\mathscr{J}_{i_{p}^{\prime}}(e)-\left\{e(1),\ldots,e\left(i_{p}^{\prime}-1\right)\right\}$ for $p\in\{1,\ldots,k\}$. Now, consider the function $\bar{e}$ such that $\bar{e}(i)=\min\left\{\mathscr{J}_{i}(e)\cap\{e(1),\ldots,e(i-1)\}\right\}$, $\forall i\in\mathscr{I}^{\prime}$, and $\bar{e}(i)=e(i)$, otherwise. So, we have $\bar{e}\in\bar{E}$, $\mathscr{I}_{j}(\bar{e})=\mathscr{I}_{j}(e)$ $\left(\forall j\in\mathscr{J}-\mathscr{J}^{\prime}\right)$ and $\mathscr{I}_{j}(\bar{e})\subseteq\mathscr{I}_{j}(e)$ $\left(\forall j\in\mathscr{J}^{\prime}\right)$. Therefore, $x(\bar{e})_{j}=x(e)_{j}$, $\forall j\in\mathscr{J}-\mathscr{J}^{\prime}$. Moreover, for each $j\in\mathscr{J}^{\prime}$ we have $x(e)_{j}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\right\}\in\left\{L_{j},U_{j}\right\}\quad$ (from Corollary 4), $x(\bar{e})_{j}=L_{j}$ if $\mathscr{I}_{j}(\bar{e})=\varnothing$ (from (15)) and $x(\bar{e})_{j}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(\bar{e})}S_{ij}^{\prime}\right\}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\right\}=x(e)_{j}$ if $\mathscr{I}_{j}(\bar{e})\neq\varnothing$ (from Remark 6). Hence, in any case, $x(\bar{e})_{j}\leq x(e)_{j}$ that implies $\sum_{j\in\mathscr{J}}c_{j}x(\bar{e})_{j}\leq\sum_{j\in\mathscr{J}}c_{j}x(e)_{j}$. ∎

### v2

@@ 3 Resolution of the feasible solution set for Problem (1)
Consider the problem stated in Example 2.17. According to Definition 3.5 and Table 4 (in Section 6), we obtain $\mathscr{J}_{1}=\{1,2,5\}$, $\mathscr{J}_{2}=\{1,8\}$, $\mathscr{J}_{3}=\{6,7,8\}$, $\mathscr{J}_{4}=\{2,3,4\}$, $\mathscr{J}_{5}=\{1,3,4,5\}$, $\mathscr{J}_{6}=\{8,9\}$, $\mathscr{J}_{7}=\{3,8,10\}$, $\mathscr{J}_{8}=\{9\}$, $\mathscr{J}_{9}=\{3,10\}$ and $\mathscr{J}_{10}=\{6,8\}$. Hence, according to Remark 3.9, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{I}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 3\times 4\times 2\times 3\times 1\times 2\times 2=5184$. Now, noting Definition 3.6 and Corollary 3.8, consider three functions $e_{1}=[5,1,6,3,5,8,10,9,10,6]$, $e_{2}=[1,8,6,2,3,8,10,9,10,6]$ and $e_{3}=[1,8,6,2,1,8,10,9,10,6]$ from $\mathscr{I}$ to $\bigcup_{i\in\mathscr{I}}\mathscr{J}_{i}$ so that $e_{p}(i)\in\mathscr{J}_{i}$ for each $i\in\mathscr{I}$ and each $p\in\{1,2,3\}$. For function $e_{1}$, we have $e_{1}(1)=e_{1}(5)=5$; $\mathscr{I}_{5}\left(e_{1},5\right)=\{1\}$ (from Part (II) of Definition 3.6); $S_{15}^{\prime}=\{0.6\}$ and $S_{55}^{\prime}=\{0.3\}$ (from Table 4). Therefore, $\mathscr{I}_{5}\left(e_{1},5\right)\neq\varnothing$ and $S_{55}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{5}\left(e_{1},5\right)}S_{k5}^{\prime}\right)=S_{55}^{\prime}\cap S_{15}^{\prime}=\varnothing$, which implies $5\notin\mathscr{J}_{5}\left(e_{1}\right)$ (Definition 3.6, Part (III)). So, since $e_{1}(5)=5\notin\mathscr{J}_{5}\left(e_{1}\right)$, from Definition 3.6 it follows that $e_{1}$ is not an admissible function. However, both $e_{2}$ and $e_{3}$ are admissible. For instance, for admissible function $e_{2}$ we have $\mathscr{I}_{1}\left(e_{2}\right)=\{1\}$, $\mathscr{I}_{2}\left(e_{2}\right)=\{4\}$, $\mathscr{I}_{3}\left(e_{2}\right)=\{5\}$, $\mathscr{I}_{4}\left(e_{2}\right)=\varnothing$, $\mathscr{J}_{5}\left(e_{2}\right)=\varnothing$, $\mathscr{I}_{6}\left(e_{2}\right)=\{3,10\}$, $\mathscr{I}_{7}\left(e_{2}\right)=\varnothing$, $\mathscr{I}_{8}\left(e_{2}\right)=\{2,6\}$, $\mathscr{I}_{9}\left(e_{2}\right)=\{8\}$ and $\mathscr{I}_{10}\left(e_{2}\right)=\{7,9\}$. So, by noting Table 4, it follows that $\bigcap_{i\in\mathscr{I}_{1}\left(e_{2}\right)}S_{i1}^{\prime}=S_{11}^{\prime}=\{0.4\}$, $\bigcap_{i\in\mathscr{I}_{2}\left(e_{2}\right)}S_{i2}^{\prime}=S_{42}^{\prime}=\{0.64\}$, $\bigcap_{i\in\mathscr{I}_{3}\left(e_{2}\right)}S_{i3}^{\prime}=S_{53}^{\prime}=\{0.6\}$, $\bigcap_{i\in\mathscr{I}_{6}\left(e_{2}\right)}S_{i6}^{\prime}=S_{36}^{\prime}\cap S_{10,6}^{\prime}=\{0,1\}$, $\bigcap_{i\in\mathscr{I}_{8}\left(e_{2}\right)}S_{i8}^{\prime}=S_{28}^{\prime}\cap S_{68}^{\prime}=\{0.2\}$, $\bigcap_{i\in\mathscr{I}_{9}\left(e_{2}\right)}S_{i9}^{\prime}=S_{89}^{\prime}=\{0.8\}$ and $\bigcap_{i\in\mathscr{I}_{10}\left(e_{2}\right)}S_{i,10}^{\prime}=S_{7,10}^{\prime}\cap S_{9,10}^{\prime}=\{0.6\}$.
@@
Furthermore, from Table 3, $I_{4}=[0,1]$, $I_{5}=[0.3,0.6]$ and $I_{7}=[0,0.55]$. Hence, from Definition 3.7, $S\left(e_{2}\right)$ is calculated as the Cartesian product $S\left(e_{2}\right)=\{0.4\}\times\{0.64\}\times\{0.6\}\times[0,1]\times[0.3,0.6]\times\{0,1\}\times[0,0.55]\times\{0.2\}\times\{0.8\}\times\{0.6\}$.
@@ 4 Optimal solutions of Problem (1)
Clearly, $S^{*}\subseteq S\left(A^{+},A^{-},b\right)$. Also, from (11) and (15), it follows that if $x\left(e_{0}\right)\in F$ (for some $e_{0}\in E$ ), then $x\left(e_{0}\right)\in S\left(e_{0}\right)\subseteq\bigcup_{e\in E}S(e)$. So, by Theorem 3.10 , it is concluded that $F\subseteq S\left(A^{+},A^{-},b\right)$. To complete the proof, it is sufficient to show that any solution $x\in S\left(A^{+},A^{-},b\right)-F$ cannot be an optimal solution. For this purpose, let $x\in S\left(A^{+},A^{-},b\right)-F$. So, from Theorem 3.10, there exists at least one $e_{0}\in E$ such that $x\in S\left(e_{0}\right)$. Now, by (11) and (15), we have $x\left(e_{0}\right)_{j}\leq x_{j}$, $\forall j\in\mathscr{J}$; and since $x\notin F$, at least one of the preceding inequalities must be strict. Without loss of generality, assume that $x\left(e_{0}\right)_{j_{0}}<x_{j_{0}}$. Hence, $c_{j_{0}}x\left(e_{0}\right)_{j_{0}}+\sum_{j\in\mathscr{J}-\left\{j_{0}\right\}}c_{j}x\left(e_{0}\right)_{j}<c_{j_{0}}x_{j_{0}}+\sum_{j\in\mathscr{J}-\left\{j_{0}\right\}}c_{j}x_{j}$, which implies $x\notin S^{*}$. ∎
Theorem 4.2 also provides a necessary optimality condition in the sense that if $x^{*}$ is an optimal solution to Problem (1), then it must belong to $F$. In other words, each solution $x(e)\in F$ is an optimal candidate solution.
@@ 5.1 Simplification techniques for finding the feasible solutions set
Under the assumptions of Lemma 5.4, the $i_{0}$’th equation is a redundant constraint. Lemma 5.4 additionally provides a further idea that is concerned with the case in which certain $x_{j}$’s can be calculated straight away without solving the problem, but just by reflecting on the special characteristics of the problem. Due to this, in such cases, any parts involved in these $x_{j}$’s can be excluded. These considerations have led us to come up with the following reduction rule.
Corollary 5.5.
@@ 5.2 Simplification techniques for finding optimal solutions
The techniques described in this subsection are based on Theorem 4.2. These techniques reduce the size of the problem in such a way that either the set $F$ does not change or the search domain is reduced to a subset of $F$ that contains optimal solutions. In fact, by using these methods, some parts of the feasible region that do not contain optimal solutions may be removed. Therefore, these methods cannot be used to find the feasible region and are only suitable for finding optimal solutions.
Lemma 5.10.
@@
It is sufficient to prove $x_{j_{0}}^{*}=L_{j_{0}}$ for any optimal solution $x^{*}$. But, since $x^{*}\in S^{*}$, Theorem 4.2 implies that $x^{*}\in F$; that is, $x^{*}=x(e)$ for some $e\in E$. Now the result follows from Lemma 5.13. ∎
Lemma 5.15.
Suppose that the simplification technique stated in Corollary 5.11 has been applied and there exist $j_{1},j_{2}\in\mathscr{J}$ such that $\mathscr{I}_{j_{1}}\neq\varnothing$ and $\mathscr{I}_{j_{1}}\subseteq\mathscr{I}_{j_{2}}$. Also, define $F_{0}=\left\{x(e):e\in E_{0}\right\}$ and $E_{0}=\left\{e\in E:\mathscr{I}_{j_{1}}(e)=\varnothing\right\}$.
@@
According to Lemma 5.15, we can reduce our search to set $F_{0}$ instead of set $F$. To achieve this, we neglect all the admissible functions $e\in E-E_{0}$ by removing each $i\in\mathscr{I}_{j_{1}}$ before selecting the vectors $e$ to construct solutions $x(e)$. However, this is equivalent to assigning $x_{j_{2}}^{*}=L_{j_{2}}$ ($x_{j_{2}}^{*}=U_{j_{2}}$ in part (b) of Lemma 5.15) and deleting the $j_{1}$’th column from the problem. The following reduction rule summarizes the previous discussion.
Corollary 5.16.
Suppose that the simplification technique stated in Corollary 5.11 has been applied.
@@ 5.3 Simplification strategies for finding optimal solutions
Definition 5.17.
Suppose that the simplification technique stated in Corollary 5.11 has been applied. An admissible function $e$ is said to be a modified function if $e(i)\in\bar{\mathscr{J}}_{i}(e)$ , $\forall i\in\mathscr{I}$, where
@@
Example 5.18.
Consider admissible functions $e_{2}=[1,8,6,2,3,8,10,9,10,6]$ and $e_{3}=[1,8,6,2,1,8,10,9,10,6]$ stated in Example 3.11. As calculated in Example 3.11, $\mathscr{J}_{5}=\{1,3,4,5\}$. Also, for admissible function $e_{2}$, we have $e_{2}(1)=1$, $e_{2}(2)=8$, $e_{2}(3)=6$ and $e_{2}(4)=2$. Furthermore, Part (II) of Definition 3.6 indicates that $\mathscr{I}_{1}\left(e_{2},5\right)=\{1\}$ and $\mathscr{I}_{3}\left(e_{2},5\right)=\mathscr{I}_{4}\left(e_{2},5\right)=\mathscr{I}_{5}\left(e_{2},5\right)=\varnothing$. On the other hand, from Table 4, it follows that $S_{51}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{1}\left(e_{2},5\right)}S_{k1}^{\prime}\right)=S_{51}^{\prime}\cap S_{11}^{\prime}=\{0.4\}$. Therefore, from Part (III) of Definition 3.6, we obtain $\mathscr{J}_{5}\left(e_{2}\right)=\{1,3,4,5\}$, which implies $\mathscr{J}_{5}\left(e_{2}\right)\cap\left\{e_{2}(1),e_{2}(2),e_{2}(3),e_{2}(4)\right\}=\{1,3,4,5\}\cap\{1,8,6,2\}=\{1\}\neq\varnothing$. So, $\bar{\mathscr{J}}_{5}\left(e_{2}\right)=\{1\}\quad$ and $e_{2}(5)=3\notin\bar{\mathscr{J}}_{5}\left(e_{2}\right)$, that means $e_{2}$ is not a modified function (Definition 5.17). However, function $e_{3}$ is an example of an admissible function that is also a modified function.
Theorem 5.19.
Suppose that the simplification technique stated in Corollary 5.11 has been applied. If $S\left(A^{+},A^{-},b\right)\neq\varnothing$ and $S^{*}$ denotes the set of optimal solutions for Problem (1), then $S^{*}\subseteq\bar{F}\subseteq F$.
@@
From Definition 5.17, it is clear that $\bar{F}\subseteq F$. To complete the proof, we shall show that for each $e\in E-\bar{E}$, there exists some $\bar{e}\in\bar{E}$ such that $\sum_{j\in\mathscr{J}}c_{j}x(\bar{e})_{j}\leq\sum_{j\in\mathscr{J}}c_{j}x(e)_{j}$. For this purpose, let $e\in E-\bar{E}$, $\mathscr{I}^{\prime}=\left\{i_{1}^{\prime},\ldots,i_{k}^{\prime}\right\}$ and $\mathscr{J}^{\prime}=\left\{j_{1}^{\prime},\ldots,j_{k}^{\prime}\right\}$ where $\mathscr{J}_{i_{p}^{\prime}}(e)\bigcap\left\{e(1),\ldots,e\left(i_{p}^{\prime}-1\right)\right\}\neq\varnothing$ and $e(i_{p}^{\prime})=j_{p}^{\prime}\in\mathscr{J}_{i_{p}^{\prime}}(e)-\left\{e(1),\ldots,e\left(i_{p}^{\prime}-1\right)\right\}$ for $p\in\{1,\ldots,k\}$. Now, consider the function $\bar{e}$ such that $\bar{e}(i)=\min\left\{\mathscr{J}_{i}(e)\cap\{e(1),\ldots,e(i-1)\}\right\}$, $\forall i\in\mathscr{I}^{\prime}$, and $\bar{e}(i)=e(i)$, otherwise. So, we have $\bar{e}\in\bar{E}$, $\mathscr{I}_{j}(\bar{e})=\mathscr{I}_{j}(e)$ $\left(\forall j\in\mathscr{J}-\mathscr{J}^{\prime}\right)$ and $\mathscr{I}_{j}(\bar{e})\subseteq\mathscr{I}_{j}(e)$ $\left(\forall j\in\mathscr{J}^{\prime}\right)$. Therefore, $x(\bar{e})_{j}=x(e)_{j}$, $\forall j\in\mathscr{J}-\mathscr{J}^{\prime}$. Moreover, for each $j\in\mathscr{J}^{\prime}$ we have $x(e)_{j}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\right\}\in\left\{L_{j},U_{j}\right\}\quad$ (from Corollary 2.16), $x(\bar{e})_{j}=L_{j}$ if $\mathscr{I}_{j}(\bar{e})=\varnothing$ (from (12)) and $x(\bar{e})_{j}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(\bar{e})}S_{ij}^{\prime}\right\}=\min\left\{\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\right\}=x(e)_{j}$ if $\mathscr{I}_{j}(\bar{e})\neq\varnothing$ (from Remark 5.12). Hence, in any case, $x(\bar{e})_{j}\leq x(e)_{j}$ that implies $\sum_{j\in\mathscr{J}}c_{j}x(\bar{e})_{j}\leq\sum_{j\in\mathscr{J}}c_{j}x(e)_{j}$. ∎